# Neural Post-Einsteinian Framework for Efficient Theory-Agnostic Tests of General Relativity with Gravitational Waves

### v1

@@ V Conclusions
In this work, we introduced the npE framework for efficient theory-agnostic GW tests of GR, using a deep-learning parameterization of non-GR theories extended from the ppE framework. The key component of the npE framework is a VAE. When trained with waveforms from a list of ppE theories, the VAE encoder finds a continuous representation for these waveforms in a dimensionally-reduced latent space. Then, the VAE decoder reconstructs similar waveforms using points in the latent space as parameters. For GW test of GR, the decoder can be used as a waveform model, with the latent space providing the non-GR parameters. Because the latent space unifies many non-GR theories that fit within the ppE framework, one only needs to run parameter estimation once to test all of them, which makes the npE framework more efficient than the ppE one. Moreover, because the gaps between ppE theories in the latent space are filled continuously with non-ppE modifications, the npE test of GR is more theory-agnostic.
In order to test the npE framework and reach the above conclusions, we focused on a parametrized modification to the inspiral GW phase of the IMRPhenomD waveform, and created a modified gravity data set with $\Phi_{\mathrm{ppE}}$ in a variety of ppE theories and from a population of BBHs. We equipped the npE framework with a two-dimensional latent space $z_{1,2}$, in which GR sits at the center ($z_{i}=0$) and the npE phase modification $\Phi_{\mathrm{npE}}$ is proportional to the polar radius $\lVert z_{i}\rVert$. We decomposed $\Phi_{\mathrm{npE}}$ as the product of $\lVert z_{i}\rVert$, a shape function $S$, and a scale function $T$. We trained the VAE network to find a representation for $S$ with the polar angle in the latent space. After training, we found that ppE theories map to radial lines that are separated and ordered in the latent space. Gaps between those ppE lines continuously extend the ppE parameterization. We then trained a secondary network to emulate $T$, so that $\Phi_{\mathrm{npE}}$ maps the unit circle $\lVert z_{i}\rVert=1$ to modifications that are comparable with GR terms, regardless of the BBH source properties. For parametrized tests of GR with GW data, we applied $\Phi_{\mathrm{npE}}$ as a phase modification to the IMRPhenomD model, and ran Bayesian parameter estimation to recover the latent parameters $z_{i}$. The prior for $z_{i}$ was chosen to be flat within $\lVert z_{i}\rVert=1$, which we referred to as the EFT prior boundary. A deviation from GR is then detected if the $z_{i}$ posterior excludes $z_{i}=0$, and this posterior then can identify and resolve the type of non-GR modification that is best supported by the data in the latent space. We note that our VAE is customized with additional features added to the vanilla formulation, including the separation of the latent angle and radius, the symmetrization of the latent representation, and the pseudo-PN expansion embedded in the decoder to better fit the physical mission.
@@
Our results have various direct applications to GW tests of GR in the upcoming LVK O5 run, and perhaps, even for O1-O4 observations. The efficiency of npE tests will be especially desirable, once the number of signals increases as expected in O5 and with next-generation detectors. For signals detected at potentially higher SNRs, the npE framework offers less biased recovery, assuming that GR is tested against modifications that deviate slightly from the standard ppE form. In addition, the npE posterior can also raise an alarm when an observed non-GR effect cannot be modelled by a PN expansion.
@@
In [48], the authors worked on an upgraded ppE framework with higher-PN-order terms, and proposed a way to regularize the prior for those higher-PN-order coefficients in order to obtain informative posteriors. This results in a more theory-agnostic version of the ppE framework that can be used for practical purposes. In comparison, our npE framework can also recover signals with higher-PN-order corrections at gaps between ppE lines in the latent space, and the EFT prior boundary serves similarly as the regularization on the prior. Apart from that, there are also regions in the latent space that are far from any ppE lines, allowing the npE test to raise an alarm when the GW signal contains a non-GR effect that cannot be modelled by any PN expansion. This features the flexibility of the neural networks, which is new in the npE framework.
Our npE framework is built on a training set composed of ppE modifications to the GW phase of the inspiral of BBHs. This training set can be augmented to suit the npE framework for more general purposes. For example, NSs can be included in the training set, provided that $\bar{f}_{\mathrm{min}}$ of the frequency grid is reduced accordingly to cover $f_{\mathrm{low}}$ of the detector with NS masses. One may also teach the VAE to learn beyond-ppE theories or modifications to the merger and the ringdown, possibly with amplitude modifications included as well. Doing so may require the shape model to be redesigned, and the dimension of the latent space to be increased.
@@ Appendix D Detecting Leading-PN-Order Deviations not Covered by the npE Training Set
Figure 15 shows examples using the npE framework to recover injections with $-1.5$PN ($b_{\mathrm{ppE,inj}}=-8$) and $0.5$PN ($b_{\mathrm{ppE,inj}}=-4$) modifications. These choices allow us to cover modifications at both negative and positive PN orders. For each case, we choose $\beta_{\mathrm{ppE,inj}}$ in reference to the boundaries shown in Fig. 9 and check whether a detection or resolution is achieved as expected. For the $-1.5$PN modification, we choose $\beta_{\mathrm{ppE,inj}}$ to be right above the interpolated resolution boundary, and we confirm that the resolution is achieved as the $z_{b}$ posterior exludes zero, and the $\varphi$ posterior excludes both neighboring ppE lines. For the $0.5$PN modification, the resolution boundary is undetermined by Fig. 9, so we choose $\beta_{\mathrm{ppE,inj}}/\beta_{\mathrm{ppE,max}}=100\%$ at maximum, which is at least above the interpolated detection boundary. In this case, we find that detection is indeed successful, as the $z_{b}$ posterior excludes zero. Although a resolution of the theory type is not expected, the npE posterior at least covers the $0.5$PN line in both the $z_{1}$-$z_{2}$ parametrization and the $z_{b}$-$\varphi$ parametrization. These observations suggest that the npE framework performs well when recovering those ppE modifications with $b_{\mathrm{ppE}}$’s not included in the training set, which proves the robustness of our npE prescription for generalizing the ppE framework.

### v2

@@ V Conclusions
In this work, we introduced the npE framework for efficient theory-agnostic GW tests of GR, using a deep-learning parametrization of non-GR theories extended from the ppE framework. The key component of the npE framework is a VAE. When trained with waveforms from a list of ppE theories, the VAE encoder finds a continuous representation for these waveforms in a dimensionally-reduced latent space. Then, the VAE decoder reconstructs similar waveforms using points in the latent space as parameters. For GW test of GR, the decoder can be used as a waveform model, with the latent space providing the non-GR parameters. Because the latent space unifies many non-GR theories that fit within the ppE framework, one only needs to run parameter estimation once to test all of them, which makes the npE framework more efficient than the ppE one. Moreover, because the gaps between ppE theories in the latent space are filled continuously with non-ppE modifications, the npE test of GR is more theory-agnostic.
In order to test the npE framework and reach the above conclusions, we focused on a parametrized modification to the inspiral GW phase of the IMRPhenomD waveform, and created a modified gravity dataset with $\Phi_{\mathrm{ppE}}$ in a variety of ppE theories and from a population of BBHs. We equipped the npE framework with a two-dimensional latent space $z_{1,2}$, in which GR sits at the center ($z_{i}=0$) and the npE phase modification $\Phi_{\mathrm{npE}}$ is proportional to the polar radius $\lVert z_{i}\rVert$. We decomposed $\Phi_{\mathrm{npE}}$ as the product of $\lVert z_{i}\rVert$, a shape function $S$, and a scale function $T$. We trained the VAE network to find a representation for $S$ with the polar angle in the latent space. After training, we found that ppE theories map to radial lines that are separated and ordered in the latent space. Gaps between those ppE lines continuously extend the ppE parametrization. We then trained a secondary network to emulate $T$, so that $\Phi_{\mathrm{npE}}$ maps the unit circle $\lVert z_{i}\rVert=1$ to modifications that are comparable with GR terms, regardless of the BBH source properties. For parametrized tests of GR with GW data, we applied $\Phi_{\mathrm{npE}}$ as a phase modification to the IMRPhenomD model, and ran Bayesian parameter estimation to recover the latent parameters $z_{i}$. The prior for $z_{i}$ was chosen to be flat within $\lVert z_{i}\rVert=1$, which we referred to as the EFT prior boundary. A deviation from GR is then detected if the $z_{i}$ posterior excludes $z_{i}=0$, and this posterior then can identify and resolve the type of non-GR modification that is best supported by the data in the latent space. We note that our VAE is customized with additional features added to the vanilla formulation, including the separation of the latent angle and radius, the symmetrization of the latent representation, and the pseudo-PN expansion embedded in the decoder to better fit the physical mission.
@@
Our results have various direct applications to GW tests of GR in the upcoming LVK O5 run, and perhaps, even for O1-O4 observations. The efficiency of npE tests will be especially desirable, once the number of signals increases as expected in O5 and with next-generation detectors. For signals detected at potentially higher SNRs, the npE framework offers less biased recovery, assuming that GR is tested against modifications that deviate slightly from the standard ppE form. In addition, the npE posterior can also raise an alarm when an observed non-GR effect cannot be modeled by a PN expansion.
@@
In [50], the authors worked on an upgraded ppE framework with higher-PN-order terms, and proposed a way to regularize the prior for those higher-PN-order coefficients in order to obtain informative posteriors. This results in a more theory-agnostic version of the ppE framework that can be used for practical purposes. In comparison, our npE framework can also recover signals with higher-PN-order corrections at gaps between ppE lines in the latent space, and the EFT prior boundary serves similarly as the regularization on the prior. Apart from that, there are also regions in the latent space that are far from any ppE lines, allowing the npE test to raise an alarm when the GW signal contains a non-GR effect that cannot be modeled by any PN expansion. This features the flexibility of the neural networks, which is new in the npE framework.
Our npE framework is built on a training set composed of ppE modifications to the GW phase of the inspiral of BBHs. This training set can be augmented to suit the npE framework for more general purposes. For example, NSs can be included in the training set, provided that $\bar{f}_{\mathrm{min}}$ of the frequency grid is reduced accordingly to cover $f_{\mathrm{low}}$ of the detector with NS masses. One may also teach the VAE to learn beyond-ppE theories or modifications to the merger and the ringdown, possibly with amplitude modifications included as well. Doing so may require the shape model to be redesigned, and the dimension of the latent space to be increased. With additional effects properly modeled beyond the inspiral phase, the boundaries in Fig. 9 can be lowered for detection of smaller non-GR deviations.
@@ Appendix D Detecting Leading-PN-Order Deviations not Covered by the npE Training Set
Figure 15 shows examples using the npE framework to recover injections with $-1.5$PN ($b_{\mathrm{ppE,inj}}=-8$) and $0.5$PN ($b_{\mathrm{ppE,inj}}=-4$) modifications. These choices allow us to cover modifications at both negative and positive PN orders. For each case, we choose $\beta_{\mathrm{ppE,inj}}$ in reference to the boundaries shown in Fig. 9 and check whether a detection or resolution is achieved as expected. For the $-1.5$PN modification, we choose $\beta_{\mathrm{ppE,inj}}$ to be right above the interpolated resolution boundary, and we confirm that the resolution is achieved as the $z_{b}$ posterior excludes zero, and the $\varphi$ posterior excludes both neighboring ppE lines. For the $0.5$PN modification, the resolution boundary is undetermined by Fig. 9, so we choose $\beta_{\mathrm{ppE,inj}}/\beta_{\mathrm{ppE,max}}=100\%$ at maximum, which is at least above the interpolated detection boundary. In this case, we find that detection is indeed successful, as the $z_{b}$ posterior excludes zero. Although a resolution of the theory type is not expected, the npE posterior at least covers the $0.5$PN line in both the $z_{1}$-$z_{2}$ parametrization and the $z_{b}$-$\varphi$ parametrization. These observations suggest that the npE framework performs well when recovering those ppE modifications with $b_{\mathrm{ppE}}$’s not included in the training set, which proves the robustness of our npE prescription for generalizing the ppE framework.